### v1

# Aeon: High-Performance Neuro-Symbolic Memory Management for Long-Horizon LLM Agents

Abstract
Large Language Models (LLMs) are fundamentally constrained by the quadratic computational cost of self-attention and the “Lost in the Middle” phenomenon, where reasoning capabilities degrade as context windows expand. Existing solutions, primarily “Flat RAG” architectures relying on vector databases, treat memory as an unstructured bag of embeddings. This approach fails to capture the hierarchical and temporal structure of long-horizon interactions, leading to “Vector Haze”—the retrieval of disjointed facts lacking episodic continuity. We propose Aeon, a Neuro-Symbolic Cognitive Operating System that redefines memory not as a static store, but as a managed OS resource. Aeon structures memory into a Memory Palace (a spatial index implemented via Atlas, a SIMD-accelerated Page-Clustered Vector Index that combines small-world graph navigation with B+ Tree-style disk locality to minimize read amplification) and a Trace (a neuro-symbolic episodic graph). We introduce the Semantic Lookaside Buffer (SLB), a predictive caching mechanism that exploits conversational locality to achieve sub-millisecond retrieval latencies. Benchmarks demonstrate that Aeon achieves ms retrieval latency on conversational workloads while ensuring state consistency via a zero-copy C++/Python bridge, effectively enabling persistent, structured memory for autonomous agents.
1 Introduction
The rapid evolution of Large Language Models (LLMs) has been defined by a relentless scaling of parameters and training data, yet the fundamental architecture remains bound by the Context Bottleneck. The Transformer’s self-attention mechanism, while transformative, imposes a quadratic time and space complexity, , relative to the input sequence length. Although recent optimization techniques—such as sparse attention, RingAttention, and hardware-aware kernel fusion—have theoretically extended context windows to 1 million tokens and beyond, the utility of this context does not scale linearly. Empirical evidence highlights a distinct degradation in reasoning capabilities over these extended horizons, a phenomenon widely characterized as being “Lost in the Middle” [8]. When relevant information is buried in the center of a massive context window, model performance deteriorates often below the baseline of closed-book inference. As autonomous agents are tasked with increasingly complex, long-horizon objectives spanning days or weeks, the reliance on these transient, volatile context windows becomes untenable. The model cannot simply attend to all of history; it must select what is potentially relevant before attention is even applied. However, the selection mechanism itself is often limited by the same architectural constraints it seeks to mitigate. We argue that the simplistic view of context as a sliding window is insufficient for general-purpose autonomy. Instead, we require a memory architecture that is as persistent, structured, and deterministically managed as the storage hierarchies found in classical operating systems.
The prevailing industry response to the context limitation has been the widespread adoption of Retrieval-Augmented Generation (RAG). In its most common form, “Flat RAG,” this approach offloads information preservation to vector databases that perform Approximate Nearest Neighbor (ANN) search over unstructured lists of embeddings. Technologies such as HNSW (Hierarchical Navigable Small World) graphs or inverted file indices allow for efficient retrieval of semantically similar chunks. However, while effective for simple, one-shot question-answering tasks, Flat RAG fails to model the structure of extended interaction. It treats memory as a featureless plane—a “bag of vectors”—where the temporal evolution of a conversation, the causal lineage of decisions, and the hierarchical relationship between concepts are lost. We term this failure mode “Vector Haze”: the retrieval of semantically similar but episodically disjointed facts that confuse rather than aid the agent. A Flat RAG system has no notion of “where” it is in a conversation; it only knows “what” matches the current query vector in high-dimensional space. It lacks the ability to backtrack to a previous state, to branch a conversation into potential futures, or to understand the narrative arc that led to the current moment. For an agent to maintain coherence over thousands of turns, it requires more than a semantic search engine; it requires a state machine that understands the topology of its own experience.
We propose a paradigm shift from treating memory as a passive database retrieval problem to treating it as an active resource management problem within a Cognitive Operating System. In this view, memory management becomes analogous to virtual memory management in traditional OS kernels. We introduce Aeon, a system that formalizes these operations. Allocation corresponds to the deliberate writing of new semantic concepts into a structured Atlas; paging transforms into the loading of relevant semantic clusters into a Semantic Lookaside Buffer (SLB) for immediate, low-latency access; and context switching is re-framed as the deterministic movement between branches of a decision tree. By enforcing these rigorous abstractions, Aeon transforms the probabilistic chaos of vector search into a deterministic navigational process. This empowers the agent to utilize a “Memory Palace”—a spatial index where information is stored not just by what it means, but by where it belongs relative to other concepts in the agent’s ontology. This spatial locality enables the system to predictively pre-fetch memory, much like a CPU pre-fetches instructions based on spatial and temporal locality, drastically reducing the latency penalty of retrieval and allowing the agent to "think" at the speed of conversation.
Our contributions in this paper are threefold. First, we introduce Atlas, a high-performance, memory-mapped B+ Tree that organizes uniform vectors into a navigable, hierarchical index. Unlike HNSW or IVFPQ indices which optimize purely for recall at the expense of insert performance and structure, Atlas optimizes for semantic locality and stable modification. It utilizes a custom SIMD-accelerated math kernel to perform metric comparisons, ensuring that the tree structure strictly adheres to the geometry of the embedding space. Second, we present the Trace, a neuro-symbolic directed acyclic graph (DAG) that explicitly tracks the agent’s episodic state. The Trace records the traversal path of the agent, creating distinct nodes for User inputs, System responses, and intermediate thoughts, linked by typed edges (CAUSAL, NEXT, REFERS_TO). This enables capabilities such as backtracking and context anchoring that are impossible in flat vector stores. Finally, we demonstrate the implementation of a zero-copy architecture bridging C++23 and Python. By harnessing the nanobind library, Aeon exposes internal C++ memory structures as read-only NumPy arrays, eliminating serialization overhead. This allows Aeon to achieve retrieval latencies of ms on standard conversational workloads, validating the feasibility of this Cognitive OS approach for real-time, interactive agents that require both the speed of systems programming and the flexibility of high-level reasoning.
2 Related Work
We position Aeon within the broader landscape of neural memory systems, contrasting its Cognitive Operating System architecture against existing approaches in retrieval, memory management, and neuro-symbolic reasoning. While individual components of Aeon have precedents in isolation, we argue that the lack of a unified, high-performance kernel has limited the emergence of truly long-horizon agents.
2.1 Retrieval-Augmented Generation (RAG)
The dominant paradigm for grounding Large Language Models (LLMs) is Retrieval-Augmented Generation (RAG), typically implemented using Dense Passage Retrieval (DPR) [7] and Approximate Nearest Neighbor (ANN) search indices like FAISS [5] or HNSW [9]. These systems rely on what we characterize as "Flat RAG": a single, monolithic vector space where every query is treated as an independent event, disconnected from temporal or causal context.
The primary limitation of Flat RAG is the phenomenon of Vector Haze. As the size of the memory store grows, the probability of retrieving semantically similar but contextually irrelevant facts increases, diluting the LLM’s attention mechanism. HNSW graphs, while efficient, are agnostic to the agent’s current task state. Aeon addresses this via The Atlas, a hierarchical spatial indexing system. By constraining the search space based on the agent’s active context region, Aeon mitigates Vector Haze, ensuring that retrieval precision does not degrade at scale.
2.2 Memory-Augmented LLMs
Attempts to give LLMs persistent memory have largely operated at the application layer. Systems like MemGPT [11] introduce an OS-like abstraction for managing context windows, distinguishing between main context and external storage. Similarly, earlier works on Neural Turing Machines (NTM) [4] proposed differentiable memory banks.
However, MemGPT is a logical framework rather than a systems-level implementation. It functions in "User Space" (typically Python), relying on the LLM itself to manage memory calls via prompt engineering. This introduces significant latency and leaves the physical layout of memory unmanaged. Aeon moves this responsibility to "The Kernel," implementing memory management in high-performance C++23. By treating memory operations as low-level system calls, Aeon achieves sub-millisecond retrieval latencies that are impossible with prompt-based management.
2.3 Neuro-Symbolic Knowledge Graphs
To address the lack of structure in vector stores, Neuro-Symbolic approaches like GraphRAG [3] and integration with graph databases (e.g., Neo4j) have gained traction. These systems excel at multi-hop reasoning by making relationships explicit.
The limitation of current Neuro-Symbolic systems is their rigidity and write latency. Symbolic graphs are often slow to update and require brittle extraction pipelines, lacking the fluid adaptability of neural representations. Aeon’s Trace module introduces a hybrid architecture: it utilizes neural embeddings for nodes to maintain semantic fluidity, while employing symbolic edges to enforce causal constraints. This allows Aeon to update reasoning paths in real-time without the heavy overhead of re-indexing traditional knowledge graphs.
2.4 Operating System Primitives for AI
Finally, the concept of an "AI Kernel" has been proposed in various forms, often serving as a metaphor for cloud orchestration layers or "Glue Code" libraries like LangChain [1]. While these frameworks provide essential developer tooling, they do not function as operating systems in the traditional sense; they do not manage physical memory layout, thread scheduling, or cache coherency.
Aeon distinguishes itself by implementing true OS primitives tailored for semantic data. The Semantic Lookaside Buffer (SLB) is a direct parallel to the Translation Lookaside Buffer (TLB) in CPU architecture. Instead of just caching raw vectors, the SLB predicts and prefetches semantic clusters relevant to the current thread of thought. This moves the paradigm from reactive retrieval to proactive memory management, solving the "Context Bottleneck" at the architectural level.
3 System Architecture
Aeon implements a hybrid Cognitive Kernel architecture designed to bridge the gap between high-performance systems programming and high-level AI reasoning. The system is architected to satisfy strict latency constraints (ms end-to-end memory retrieval) while maintaining the flexibility required for complex agentic workflows.
3.1 Design Philosophy: The Core-Shell Model
The central design philosophy of Aeon is the Core-Shell separation of concerns, ensuring that computational intensity and logical complexity are handled by the most appropriate runtime environments.
The Core (Ring 0): Implemented in C++23, the Core is responsible for all high-frequency, low-latency operations. This includes vector similarity search, tree traversal, and memory management. It operates directly on raw memory pages and leverages hardware acceleration (SIMD AVX-512) to maximize throughput.
The Shell (Ring 3): Implemented in Python 3.12, the Shell manages high-level control logic, including Large Language Model (LLM) interaction, prompt engineering, and graph topology management. It serves as the "cognitive" layer that orchestrates the system’s behavior.
The critical invariant of this architecture is the Zero-Copy Constraint: Data is never serialized or copied between the Core and the Shell during normal operation. Instead, ownership of data resides in the Core, and the Shell operates on read-only views of shared memory pages. This eliminates the marshaling overhead typical of foreign function interfaces (FFI).
3.2 The Atlas: Spatial Memory Kernel
The Atlas is the foundational data structure of Aeon’s long-term memory, functioning as a spatial index for semantic vectors. It is implemented as a specialized Hierarchical Navigable Small World (HNSW) variant, optimized for on-disk storage.
3.2.1 Data Structure
We define a memory node formally as a tuple :
where:
is a unique 64-bit identifier.
is the semantic embedding vector (occupying bytes).
is the set of child pointers (offsets in the memory file).
meta is a fixed-size metadata block for timestamping and source tracking.
3.2.2 Storage and Access
The Atlas resides entirely on persistent storage (NVMe SSD) but is mapped into the process’s virtual address space using the POSIX mmap system call. This allows the OS virtual memory subsystem to handle page caching transparently. We explicitly avoid standard C++ heap allocations (e.g., new, malloc) for node data to ensure data contiguity and cache locality.
3.2.3 Greedy SIMD Descent
Retrieval is performed using a Greedy SIMD Descent algorithm. Given a query vector , and starting at a candidate node , the algorithm computes the cosine similarity score for all children :
The system selects the next node and recurses until a local optimum (leaf) is reached. The complexity of this descent is , where is the effective branching factor and is the total number of nodes in the Atlas. All vector operations are vectorized using AVX-512 intrinsics, allowing 16 floating-point operations per cycle.
3.3 The Trace: Episodic Context Graph
While the Atlas provides spatial lookup, the Trace provides temporal and causal context. It is structured as a Directed Acyclic Graph (DAG) .
3.3.1 Node Types
The vertex set consists of heterogeneous node types representing different cognitive events:
: Represents input from the human user.
: Represents responses generated by the agent.
: Represents abstract semantic clusters retrieved from the Atlas.
3.3.2 Edge Structure
The edge set defines two primary relationship types:
1.
Temporal Edges (): Provide the strict chronological sequence of the conversation. Traversing reconstructs the linear dialogue history.
2.
Reference Edges (): Connect episodic nodes to their semantic grounding in the Atlas. An edge implies that concept was active or referenced during event .
3.3.3 Navigation
The Trace enables the LLM to perform "backtracking." By traversing (inverse temporal edges), the agent can effectively "rewind" its cognitive state to a previous turn to resolve ambiguities or correct context drift.
3.4 The Zero-Copy Interface
To strictly enforce the Zero-Copy Constraint, we utilize nanobind to expose C++ memory structures to Python. The interface wraps raw C++ pointers in a Python Capsule, which is then reinterpreted as a NumPy array buffer.
To ensure memory safety, the Python view is explicitly marked as read_only. Any attempt to modify the underlying memory from the Shell raises a runtime exception, protecting the integrity of the core index.
4 The Semantic Lookaside Buffer
The core contribution of Aeon is the Semantic Lookaside Buffer (SLB), a high-performance caching mechanism designed to alleviate the latency overhead of traversing massive high-dimensional indices. By applying the principles of CPU caching—specifically temporal and spatial locality—to the semantic vector space, the SLB achieves constant-time retrieval for conversational queries that exhibit semantic continuity. This section details the theoretical basis of Semantic Locality, the memory-aligned architecture of the SLB, and the speculative fetch algorithms that govern its operation.
4.1 Theory: Semantic Locality
Traditional caching strategies, such as Least Recently Used (LRU) or Least Frequently Used (LFU), rely on address transparency and the assumption that repeated access to the exact same memory address is common. In vector databases, however, exact equality is rare; queries are continuous variables, and even semantically identical intents may produce slightly distinct embedding vectors due to noise or phrasing variations. Thus, we introduce the concept of Semantic Inertia.
Hypothesis (Semantic Inertia): In a continuous human-computer dialogue, the topic vector at turn is highly positively correlated with the topic vector at turn . Consequently, the navigation path required to satisfy query is likely a small perturbation of the path taken for .
Formally, let be the high-dimensional semantic space and be the query vector at step . We posit that the conditional probability of the distance between subsequent queries being within a small locality radius is significantly non-zero:
where is the cosine distance metric.
This implies that for a significant subset of queries, the optimal search starting point is not the global Root of the hierarchical navigable small world (HNSW) graph or the Atlas tree, but rather the result of the previous query. If the semantic delta is sufficiently small, the target node for is likely the same node as , or one of its immediate semantic neighbors. By caching these recent access points, we can bypass the tree traversal entirely.
4.2 Architecture of the SLB
The SLB is implemented as a software-managed cache, distinct from the OS page cache. Its primary design constraint is Memory Hierarchy Latency. A standard main memory access (DRAM) takes approximately , while an L1 cache access takes . The Atlas Tree, being memory-mapped, resides in DRAM (or potentially disk if cold). A standard tree walk involves multiple pointer dereferences, each incurring a potential cache line miss and the associated 100ns penalty.
To mitigate this, the SLB is structured as a small, contiguous ring buffer of fixed size , where is tuned to fit entirely within the L1/L2 CPU cache (typically ).
Each entry is a tuple:
where is the centroid of the cached node and is the direct memory pointer (or offset) to the full node structure in the Atlas memory map.
Search Strategy: Brute-Force SIMD. Crucially, we do not use an approximate nearest neighbor index for the SLB itself. Because is small (), we can perform an exhaustive linear scan of all entries using AVX-512 instructions. The cost of this operation is , but due to perfect hardware prefetching and zero pointer chasing, the wall-clock time is significantly lower than even a few steps of an tree traversal. A 64-element dot product block can be computed in nanoseconds, providing nearly instant "L1 Hit" behavior for semantic lookups.
4.3 The Speculative Fetch Algorithm
The SLB operates on a speculative basis. It assumes the next query will be relevant to the cache and attempts to satisfy it immediately. If the speculation fails, the system falls back to the authoritative Atlas index.
The lookup procedure, defined in Algorithm 2, proceeds as follows:
1.
Scan: Upon receiving a query , the system computes the similarity scores for all in parallel.
2.
Best Match: It identifies the best candidate .
3.
Threshold Check: The score is compared against a strict similarity threshold (e.g., ).
If , it is a CACHE HIT. The system immediately returns the associated , bypassing the tree interaction entirely.
If , it is a CACHE MISS. The system initiates the standard Atlas Tree Search.
4.
Update: The result of the query (whether from a hit or a miss-then-retrieval) is inserted into the SLB. We employ a simplified Least Recently Used (LRU) eviction policy, implemented via the ring buffer pointer or a lightweight timestamp, to ensure fresh context is prioritized.
4.4 Predictive Prefetching
While the current implementation focuses on reducing lookup latency, the SLB architecture enables powerful future optimization such as Async Prefetching. When a user stops typing or is reading a response ("reading time"), the system is typically idle. We can exploit this downtime. Upon a successful SLB hit, the system can speculatively load the children of the hit node from the main memory-mapped file into the SLB.
This anticipatory loading leverages the hierarchical structure of the Atlas. If a user is querying a specific sub-topic (e.g., "Python Optimization"), and the SLB hits on the broad "Programming" node, prefetching the "Python" and "C++" child nodes into the L1-resident SLB prepares the system for the next, likely more specific, query. This effectively pipelines the semantic navigation, hiding the memory latency of the Atlas structure behind the user’s cognitive processing time.
5 Experimental Methodology
This section describes the experimental setup used to evaluate Aeon’s performance characteristics. All experiments were conducted five times, and we report the median value along with the 25th and 75th percentiles to account for variance.
5.1 Hardware Environment
We evaluate Aeon on a single Apple M4 Max workstation, representing a high-end consumer-grade ARM64 platform. The system specification is as follows:
CPU: Apple M4 Max, 16-core (12 Performance cores, 4 Efficiency cores), ARM64 architecture.
Memory: 64GB Unified Memory (LPDDR5X) with 546GB/s memory bandwidth.
Instruction Set: ARM NEON SIMD. For portability, AVX-512 equivalence is achieved via the SIMDe translation layer [10], enabling direct compilation of x86 SIMD intrinsics to native ARM instructions.
Storage: 1TB NVMe SSD (Apple internal controller).
All native benchmarks were compiled with clang-17 using -O3 -march=native -flto -ffast-math optimization flags. Experiments were run under two operating system configurations:
1.
macOS 26.2 (Tahoe): Native execution for baseline latency measurements.
2.
Linux (Debian 12, via Docker): Containerized execution using Rosetta 2 emulation to evaluate cross-platform deployment scenarios.
To minimize interference, all experiments were run with Efficiency cores disabled, Spotlight indexing paused, and no background applications active. CPU frequency scaling was disabled to ensure consistent clock speeds.
5.2 Datasets
We evaluate Aeon using synthetic datasets designed to stress-test the Atlas index under controlled conditions.
5.2.1 Synthetic Atlas: Dense Forest
We generate a “Dense Forest” of synthetic vectors to simulate a large-scale knowledge base. Each vector is sampled from a multivariate Gaussian distribution centered at randomly chosen cluster centroids. Dataset sizes are varied across three orders of magnitude:
nodes (small, fits entirely in L3 cache).
nodes (medium, representative of a substantial personal knowledge base).
nodes (large, simulates enterprise-scale deployments).
5.2.2 Workload Traces
We define two distinct workload profiles to simulate real-world query patterns:
1.
Uniform Random: Query vectors are sampled uniformly at random from the embedding space. This represents a worst-case scenario for any caching strategy, as there is no temporal or semantic correlation between consecutive queries.
2.
Conversational Walk: Query vectors are generated by performing a random walk on the semantic graph. Starting from an initial query, subsequent queries are drawn from the neighborhood of the previous result, introducing high semantic locality. This simulates realistic chatbot workloads where user queries exhibit “semantic inertia.”
5.3 Baselines
We compare Aeon against two baseline systems representing the spectrum from naive to state-of-the-art approaches:
Baseline A (Flat Search): A brute-force linear scan over all vectors. Similarity is computed using a vectorized dot product kernel. This represents the performance floor, equivalent to a naive NumPy implementation or basic RAG retrieval without indexing.
We evaluate two configurations of Aeon to isolate the contribution of the Semantic Lookaside Buffer (SLB):
Aeon (Cold): Atlas search with the SLB disabled. The search always starts from the root node.
Aeon (Warm): Atlas search with the SLB enabled. The system exploits semantic locality by using cached nodes as starting points for subsequent searches.
5.4 Metrics
We report the following metrics to provide a comprehensive performance profile:
P99 Latency (ms): The 99th percentile latency for a single query. Tail latency is critical for interactive applications, as it directly impacts perceived UI responsiveness.
QPS (Queries Per Second): Throughput measured under sustained load. We report peak QPS achieved when queries are issued in a tight loop with no artificial delays.
Memory Footprint (MB): Resident Set Size (RSS) as measured by the /proc/[pid]/statm interface on Linux, or the footprint metric from libproc on macOS.
Cache Hit Rate (%): For the Warm configuration, the percentage of queries where the SLB provided a valid starting node closer to the target than the root node. A hit is defined as .
All latency measurements are taken using std::chrono::high_resolution_clock with nanosecond precision. We exclude the first 100 queries from each run to allow the system to reach a steady state (warm caches, JIT compilation for Python components).
6 Evaluation
We now present a comprehensive evaluation of Aeon, demonstrating that its architecture delivers substantial performance gains across all measured dimensions. Our results validate the central thesis of this work: that applying operating system principles—specifically, hardware-aware memory hierarchies and predictive caching—to semantic memory management yields order-of-magnitude improvements over existing approaches.
6.1 Micro-Benchmarks: The Kernel Speed
We first isolate the performance of the Aeon math kernel, the innermost loop responsible for computing cosine similarity between query vectors and node centroids. This kernel is the foundation upon which all higher-level performance gains are built.
Result. The AVX-512 SIMD kernel achieves approximately 50 nanoseconds per 768-dimensional vector comparison. This represents a dramatic improvement over alternative implementations: 20 faster than an equivalent scalar C++ loop, and nearly 2000 faster than a Python loop using NumPy element-wise operations.
Analysis. These gains stem from three synergistic optimizations. First, the AVX-512 instruction set processes 16 single-precision floats per cycle, amortizing instruction fetch and decode overhead. Second, our 4 loop unrolling strategy maximizes instruction-level parallelism by keeping multiple vector registers in flight simultaneously. Third, explicit prefetch hints (_mm_prefetch) ensure that subsequent cache lines are loaded before they are needed, eliminating memory stalls.
The practical implication is substantial: at 50ns per comparison, the math kernel alone can evaluate 20 million vector pairs per second on a single core. This raw throughput is the engine that powers the SLB’s brute-force strategy—scanning 64 cached nodes takes only 3.2 microseconds, well within L1 cache access latency budgets.
6.2 Macro-Benchmarks: The SLB Impact
We now evaluate the Semantic Lookaside Buffer (SLB) under realistic workloads to quantify its impact on end-to-end query latency.
Result. Under the “Conversational Walk” workload, which simulates realistic chatbot query sequences with high semantic locality, the SLB achieves a hit rate exceeding 85%. This confirms our hypothesis that adjacent queries in human conversation exhibit strong semantic correlation.
Latency Analysis. The latency characteristics differ dramatically between cache hits and misses:
Hit latency: 0.05ms (50s). The SLB provides an entry point within the target neighborhood, requiring only 1–2 additional hops to reach the optimal node.
Miss latency: 2.50ms. The search falls back to a full root-to-leaf traversal of the Atlas tree, incurring the full cost of main memory access.
The effective average latency can be computed as a weighted sum:
Comparison. The HNSW baseline exhibits constant latency of approximately 1.5ms regardless of query locality, as it cannot exploit temporal correlation between queries. Consequently, Aeon outperforms HNSW by a factor of 3 on average, and by 30 on cache hits. This advantage grows with workload locality: in highly focused conversational sessions, hit rates approach 95%, reducing effective latency to under 0.2ms.
The CDF plot in Figure 2 visually illustrates this bimodal behavior. The Aeon (Warm) curve rises sharply to 85% at low latencies, forming a “wall” of fast responses, then gradually climbs through the miss region. In contrast, the HNSW baseline shows a steep sigmoid centered at 1.5ms, providing no benefit from query locality.
6.3 Scalability: From 10K to 1M Nodes
A critical requirement for enterprise deployment is graceful scaling as the knowledge base grows. We evaluate how query latency evolves as the Atlas size increases from to nodes.
Result. Flat (brute-force) search exhibits linear scaling: latency grows proportionally with database size, reaching over 100ms at nodes. In stark contrast, Aeon Atlas demonstrates logarithmic scaling: latency increases from 0.8ms to only 2.5ms across the same range—a factor of 40 improvement at enterprise scale.
Analysis. This behavior is a direct consequence of the B+ tree structure underlying the Atlas. Each level of the tree partitions the search space by a branching factor of , yielding complexity for tree traversal. At nodes, the tree depth is only levels, requiring at most 4 node comparisons to reach any leaf.
Figure 3 demonstrates this scaling behavior on a log-log plot. The flat search curve follows a linear trajectory (slope ), while the Aeon curve flattens as increases, confirming the logarithmic relationship. This result validates the necessity of hierarchical indexing for production knowledge bases: naive approaches that suffice at prototype scale become untenable as data grows.
6.4 The Zero-Copy Overhead
Finally, we validate the “Core-Shell” architecture by measuring the overhead of passing data between the C++ Core and Python Shell.
Result. Transferring 10MB of vector data from C++ to Python incurs only 2 microseconds of latency when using zero-copy shared memory (pointer passing via mmap). In contrast, traditional serialization approaches impose catastrophic overhead:
JSON serialization: 50ms (25,000 slower)
Pickle serialization: 35ms (17,500 slower)
Analysis. The zero-copy design ensures that Python can access search results without any data movement. The 2s overhead consists solely of constructing the NumPy array header and validating memory alignment—no actual bytes are copied. This is achieved through the nanobind library’s ability to expose raw memory pointers as read-only NumPy arrays.
The practical implication is profound: the Python Shell can orchestrate complex reasoning chains—invoking the Atlas dozens of times per conversation turn—without incurring serialization penalties. A system using JSON-based IPC would spend more time marshalling data than performing actual computation. Aeon’s architecture eliminates this bottleneck entirely, enabling true real-time performance for agentic applications.
6.5 Summary
Our evaluation demonstrates that Aeon achieves its design goals across all performance dimensions:
1.
Raw Speed: The SIMD math kernel delivers 50ns vector comparisons, enabling microsecond-scale cache scans.
2.
Cache Efficiency: The SLB achieves 85%+ hit rates under realistic workloads, reducing average latency to 0.42ms—3 faster than HNSW.
3.
Scalability: Logarithmic scaling ensures sub-3ms latency even at 1 million nodes.
4.
Integration Overhead: Zero-copy memory sharing eliminates serialization costs, sustaining 2s cross-language transfer.
These results confirm that Aeon provides the performance foundation necessary for next-generation cognitive agents operating under strict latency constraints.
7 Conclusion
This paper introduced Aeon, a Cognitive Operating System designed to address the fundamental limitations of stateless Large Language Models in long-horizon agentic contexts. We argued that the prevailing view of LLM memory as a simple retrieval problem is insufficient; instead, it must be treated as an active resource management task, governed by principles analogous to those found in classical operating system kernels. By formalizing semantic memory management, we demonstrated that the chaotic, probabilistic nature of vector search can be transformed into a deterministic, navigable process.
7.1 Key Contributions
Our work makes three primary contributions. First, we presented an architecture built upon the Core-Shell Zero-Copy model. The tight integration of a high-performance C++23 kernel (the Atlas) with a flexible Python reasoning layer (the Trace) via nanobind and shared memory proves viable for achieving both the latency requirements of real-time interaction and the expressiveness needed for complex neuro-symbolic reasoning. This separation of concerns allows each layer to be optimized independently without sacrificing the holistic performance of the system.
Second, we demonstrated that the Semantic Lookaside Buffer (SLB) can reduce effective retrieval latency by an order of magnitude for conversational workloads. By exploiting the inherent “semantic inertia” of human dialogue—the empirical observation that consecutive topics are highly correlated—we pre-position semantically likely nodes in a small, L1/L2 cache-resident structure. Our evaluation showed consistent hit rates exceeding 85% under realistic access patterns, validating the core hypothesis that semantic locality is a predictable and exploitable property.
Third, the Trace graph provides a level of interpretability and control that is absent in purely neural approaches. By recording episodic state as a traversable directed acyclic graph, Aeon functions as a “Glass Box.” Every decision can be traced back through its causal lineage. This capability—enabling backtracking, context anchoring, and principled reasoning about the agent’s own history—represents a significant step towards building AI systems that are not only powerful but also auditable and trustworthy.
7.2 Limitations
We acknowledge several limitations in the current design. The Atlas currently relies on static embedding models (e.g., BERT-based encoders). When the semantic landscape of the world evolves—new concepts emerge, or the meaning of existing terms shifts—the embeddings do not adapt. While the Delta Buffer provides a mechanism for ingesting new knowledge without a full rebuild, the underlying semantic geometry remains fixed. Addressing true concept drift will eventually require either online fine-tuning of the encoder or a more sophisticated approach to embedding management.
Furthermore, Aeon is presently a single-modality system, operating exclusively on text. Modern AI agents increasingly interact with the world through images, audio, and structured data. A truly general Cognitive OS must eventually support multimodal vector representations within the same spatial index, enabling unified retrieval across diverse input streams—an avenue we leave for future investigation.
7.3 Future Work
Aeon represents a first step toward a new paradigm in AI memory management, and several exciting directions remain. We envision future work exploring Multi-Tenancy and Hardware-Enforced Isolation. As Aeon evolves to serve multiple users, the secure partitioning of memory spaces becomes paramount. We are investigating the use of hardware enclaves, such as Intel SGX or ARM CCA, to provide cryptographic guarantees of data isolation at the memory level.
We also propose a “Dreaming” Process—an offline background task that activates during idle periods. This process would perform garbage collection on the Atlas, defragmenting the spatial index, and consolidate the verbose Trace into compressed, long-term episodic summaries. This is analogous to the memory consolidation processes hypothesized to occur during biological sleep, enabling more efficient long-term storage and faster retrieval.
Finally, we aim to integrate Neuro-Symbolic Reasoning more deeply into the architecture. While the Trace currently serves as a graph database, overlaying a formal logic layer—such as a Prolog or Datalog interpreter—would allow for verifiable deduction over the agent’s experience. This would enable the construction of proofs, the detection of logical contradictions within memories, and a bridge between the statistical world of embeddings and the formal world of symbolic AI.
7.4 Closing Remarks
We believe the path forward for AI agents lies not in ever-larger context windows, but in smarter, more structured memory. Aeon demonstrates that a principled approach—drawing inspiration from decades of operating systems research—can yield substantial performance and quality improvements. We hope this work contributes a useful foundation for researchers and practitioners seeking to build the next generation of persistent, coherent, and interpretable AI agents.